OpenStudy (anonymous):

What is the height of a rectangular prism with length 5 feet, width 2 feet, and volume 120 cubic feet?

jimthompson5910 (jim_thompson5910):

V = L*W*H 120 = 5*2*H solve for H

OpenStudy (anonymous):

H = 12?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Thank you for your help c:

jimthompson5910 (jim_thompson5910):

yw